A Complete Bibliography of Publications in the ICES Journal of Marine Science (1920–1929)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.04

Title word cross-reference

1929 [LA29].

Abweichungen [vR29]. account [dB27]. Action [BW27, BW29, Hod27b]. ADDENDA [Ano26a]. Aden [ME28]. adjacent [Joh28]. Age [Hod27a, Jes26, TP27, Moh27]. Aid [Ble27a]. Alters [Moh27]. Alters- [Moh27]. America [Tho29b]. amongst [LA29]. Analysis [Buc29b, HR27]. Anglian [Hod27a]. Animal [Vol28]. Apparatus [Lum27, Pet28, Sto28, Tho29a, Wol28, Hic27, Sun29]. April [LA29]. Arbeit [Sch28c]. Arcto [Sun27a]. Arcto-Norwegian [Sun27a]. Area [Pet26c, Pet26d, Sch28c, Sch28e]. assessment [Sch28b]. Atlantic [Sta27, Wat29, Le 27a, Le 29, Wat29, dB27]. Atlantique [Le 27a, Le 29]. Atlantische [Sta27]. Atlantischen [Wat29]. Augusto [Ano26h]. Autumn [Hod27a, Hes27]. autumn-spawning [Hes27].

Baltic [Hes27]. Barents [ZBI28]. basics [Ehr27a]. Bedeutung [Ehr27a].

Belt [Joh26a, Joh28]. Bestimmung [Moh27]. between [RE26]. Beurteilung [Sch28b]. Bibliographie [Moh27]. Bibliography [All26a, All26b, Ano26b, Ano26c, Ano27a, Ano27b, Ano27c, Ano28a, Ano28b, Ano28c, Ano29a, Ano29b, Ano29c, Moh27]. Biological [Büc27, Hes27, Hod27a, Sun27a, TP27, Hes27]. Biologie [Le 27a, Le 28]. Biology [Le 27a, Le 28]. Bionomics [All26a, All26b]. Bluefin [Le 27a, Le 29]. Bottles [Car28]. Bottom [ZBI28, Hic27, Wim28]. Bucht [Büc27]. Bucket [Wim28].

C. [R.28b]. Calanus [Far27]. calculating [Sun29]. Calculation [Sun26b]. carried [dB27]. catch [Ehr27a]. Caught [RE26]. Causes [Joh26a]. Channel [Le 28, Sch27a]. Charts [LC28]. Chlorine [ME28]. classes [Hjo26]. Climate [Pet26d]. Coast [Far27, Tho29b]. Cod [BW29, Dav29, RE26, Gra26a, Sun27a]. Cod-End [BW29, Dav29, RE26]. Cod-Ends [BW29]. Collection [Gra29a, Lum27, Wim28]. Colour [Sun28b]. Comparator [Sun28b]. compared [BW29]. Comparison [HR27]. Composition [Hod27a, TP27]. concerned [Atk26a, Atk26b]. concerning [Hod27b, Mol29, Pet26b]. conditions [Pal26]. Conical [Sei29]. Connection [Car28]. consideration [Atk26a, Atk26b]. content [vR29]. Contribution [Le 27a, Le 27a]. Copper [Atk27]. Correction [Sun26a]. Correspondence [Le 27b, Ort28, Pet26a, Sun27b, TE29]. Courses [Ekm28a]. Cumberland [Jes26]. Current [Ano26b, Ano26c, Ano27a, Ano27b, Ano27c, Ano28a, Ano28b, Ano28c, Ano29a, Ano29b, Ano29c, BW25, Car26, Sch27b]. Currents [Car28, Ekm28a, Ekm28b, Pet26c]. Curves [BWH29, J.26a].

Daily [Hav28]. dans [Le 27a, Le 29]. data [Sun29]. dealing [LC28]. decade [vR29]. Dekaden [vR29]. Dekaden-Abweichungen [vR29]. depth [Sun26b]. detachable [Wim28]. Determination [Buc29a, Jes26, Moh27]. Determinations [Ble27a]. Determining [Sto28, Gra26a]. Deutsche [Sta27]. Deutschen [Büc27, Wat29]. development [Ehr27a]. deviations [vR29]. Diameters [Sei29]. Different [Buc29a, Dav29]. Directions [Tån27]. Distribution [LC28, Söm29, Pal26]. Drift [BW27, Car28, Hod27b]. Durchlüftung [Wat29]. during [Hav28, Joh26a, LA29, TP27]. Dutch [Tes28]. dynamic [Sun26b].

East [Har29, Hod27a, Le 27a, Le 29]. Eastern [ME28]. Editorial [Ano26d]. Eduardo [Ano26h]. Eel [Sch27a, Tes28]. Eels [Jes26]. End [BW29, Dav29, RE26]. Ends [BW29]. English [Le 28]. Entwicklung [Ehr27a]. Erforschung [Ehr27a]. Ergebnissen [Wat29]. Est [Le 27a, Le 29]. étude [Le 27a]. Euphausiidae [Ruu28]. Expedition [Wat29, Sta27, Wat29]. Experiment [LC28]. Experimental [Dav29]. Experiments [Hod27b, Joh28, RE26]. Exploratory [Hic28].

factors [Atk26a, Atk26b]. Faroe [Sch27a]. finding [Hic27]. Finland [Pal26]. finmarchicus [Far27]. Finnischen [Pal26]. first [Gra26a]. Fischen [Moh27]. Fish [Büc27, Gra29a, Hes27, Hod27a, LA29, Pet26c, RE26, Sun27a, Tån27, Tho29a, TP27, Wol28, Moh27]. Fish-Migrations [Pet26c]. Fisheries [Pet26d, Tho29b]. Fishery [All26a, All26b, BWH29, Car26, Car28]. fishes [Hjo26]. fishing [Atk27]. Fleetwood [Hic28]. Flødevigen [DM29]. flow [Pal26]. Fluctuations [Hjo26, Ruu28, Vol28, TP27]. following [Jen27]. food [Hjo26]. Frameless [Knu29a]. French [Le 27a, Le 28, Le 29]. Frequency [BWH29, J.26a]. Freshwater [Tes28]. Friedrich [M.29].

G [R.28b]. Gadoid [Gra27a]. Gebiete [Sch28c, Sch28e]. Gelder [BW29]. Georg [H.27d]. German [Sta27, Buc29b, Ehr27a, Moh27, Pal26, Sch28b, Sch28c, Sch28e, Sta27, Wat29, vR29]. Gibraltar [Sch28c, Sch28e, dB27]. Gibraltarstrasse [Sch28c, Sch28e]. Giesbrecht [Söm29]. Graphical [Sun26b]. Growth [Far28, Hav28, Tes28, Atk26a, Atk26b, Moh27]. Growth-rate [Far28]. Grundlagen [Ehr27a]. Gulf [ME28, Pal26].

Haddock [Joh26a, Mol29]. Hake [Hic28]. Half [Sei29]. Half-Meter [Sei29]. Halibut [Tho29b]. Hansen [Ano26f]. Heincke [M.29]. Herring [Hes27, Hod27a, Hod28, J.26a, Joh27, Lea29, Ehr27a]. Herrings [Far28, Jen27]. History [Lea29]. Hydrodynamics [Har29]. Hydrography [Pet26d].

ihre [Sch28b]. II [Ano27d, Hes27, Hod27a]. III [Ano28d, TP27]. im [Büc27, Pal26, Sch28c, Sch28e, vR29]. importance [Ehr27a]. important [Hjo26]. Improvement [Lum28]. Improvised [Hic27]. Index [Ano26e, Ano27d, Ano28d, Ano29d]. Individuals [Vol28]. Influence [Jen27]. information [Wat29]. Instrument [Car26]. Interpretation [LC28]. Investigation [All26a, All26b, Car28, J.26a]. Investigations [Ekm28b, Tån27, Tes28]. Ireland [Far27, Har29]. Irish [Far28]. IV [Ano29d].

Jahrgangs [Büc27]. Jespersen [Wim28]. Joh [R.28b].

Kara [ZBI28]. Kattegat [Mol29].

L. [Le 27a, Le 29]. L’Amiral [Ano26h]. Larvae [DM29, Sch27a]. larval [Ruu28]. leading [Joh26a]. Life [Lea29]. Limfjord [Joh28]. Limit [Söm29]. living [Vol28]. Lobster [DM29].

Machine [Gra27a]. Magnesium [ME28]. Manche [Le 28]. maps [vR29]. March [LA29]. Maregraph [Sho29]. Marine [All26a, All26b, Buc29b]. Materials [ZBI28]. Measurements [Pet28]. Measuring [Car26, Tho29a, Wol28]. Mediterranean [ME28]. Meerbusen [Pal26]. Meeresforschung [Buc29b]. Mer [Le 27a, Le 29]. Mesh [RE26]. Meshes [Dav29]. Meter [BW25, Sei29]. Method [BWH29, HR27, Gra26a].

Methods [Atk27]. Migrations [Joh27, Pet26c]. Miss [Ano26f]. Mitteilung [Wat29]. Modification [Lum28]. Monats [vR29]. Monats- [vR29]. Monatskarten [vR29]. monthly [vR29]. Mortality [LA29]. Motion [Lum27]. movements [Sch28c, Sch28e].

nasutus [Söm29]. necrologiques [Ano26h]. Net [BW27, Wim28]. Nets [BW27, Hod27b, Sei29, Atk27]. Neuparth [Ano26h]. Nitrogen [Buc29b]. Nord [Le 27a, Le 29, Le 29]. Nord-Est [Le 27a]. Nordsee [vR29]. Norfolk [Jes26]. North [Hod28, Le 29, Pet26b, Tho29b, TP27, Joh28, Le 27a, Le 29, LA29, vR29]. North-East [Le 27a]. Northern [Söm29]. Norwegian [Lea29, Sun27a]. Note [Ano26d, Dav29, Ekm28a, Far28, Hod27b, Söm29, Lum27, Sch28c]. Notes [Ano26g, Le 28, Sch27b]. Notices [Ano26h]. Notiz [Sch28c]. Number [Vol28].

Oberflächensalzgehalt [vR29]. Observation [Lum28]. Observations [Le 28, Lum29, Mol29, Lum27]. Ocean [Ekm28a, Ekm28b, Wat29, dB27]. Ocean-Currents [Ekm28a, Ekm28b]. Oceanic [Lea29]. Oceanographical [Sun29]. off [Far27, Ruu28]. Oil [Rob26]. one [Sei29]. open [Sho29]. Ossian [H.27d]. Ostenfeld [Wim28]. other [BW29]. Otolith [Ble27a]. Otolithen [Sch28b]. Otoliths [Sch28b]. Oxygen [Buc29b]. Oysters [Hav28]. Ozeans [Wat29].

Pacific [Tho29b]. part [vR29]. Patterns [Sei29]. Period [TP27]. Petersen [R.28b]. pH [Buc29a]. Places [Hod28]. Plaice [Joh28, TP27]. Plankton [HR27, Pet28, Sei29, Wim28]. Plankton-Nets [Sei29]. plant [Atk26a, Atk26b]. Pollution [Rob26]. precise [Gra26a]. Preliminary [Dav29, Hod27b, RE26, Wat29]. Preparation [Gra29a]. presence [Le 27a, Le 27a]. preserve [Atk27]. Productivity [ZBI28]. Purposes [Car26].

quantitative [Atk26a, Atk26b, Pet28]. Quantities [Joh26a]. Quantity [Jen27].

Racial [Tån27]. Rate [Hav28, Far28]. Ratio [ME28]. Rearing [DM29]. Recording [BW25]. regarding [Hes27]. Regulation [Tho29b]. Relation [RE26]. Remarkable [Joh26a]. Remarks [BW27]. Reproducing [Gra27a]. Reproduction [Far27]. Research [Buc29b, Car26, Car28, Ehr27a]. results [BWH29, Wat29]. Reversing [Knu29a, Hic27]. Reviews [A.26, A.27a, A27b, And27, Ano26i, Ano26j, Ano26k, Ano26l, Ano28e, Atk26c, B.26a, B.26b, B.26c, B.26d, B.27, Bck28, Bie28, Bje28a, Bje28b, Ble27b, Ble27c, Ble28, Ble29a, Ble29b, Ble29c, C.26a, C.26b, C.26c, C.28, C.29, Car29, Dav26, E.29, Eds26, Eh.26, Eh.27, Ehr27b, F.26c, F.26d, F.26e, F.26a, F.26b, F.27a, F.27b, F.27c, F.27d, F.27e, F.28b, F.28a, F.28c, F.29a, F.29c, F.29d, F.29b, Gil26, Gil27, Gil28, Gra26b, Gra26c, Gra27b, Gra27c, Gra29b, Gra28, H.26, H.27e, H.27b, H.27c, H.27a, H28b, H.28a, H.29d, H.29c, H.29a, H.29b, H.29e, Har27, Hel26, Hod26, Hun27, J.26b, J.27, J.29a, J.29b, Jac28, Jes28a, Jes28b, Joh26b, Knu29b, Knu29c, Kün29, L.26a, L.26b, L.27a, L.27b, L.28, L.29a, L.29b, L.29c, M.26, M.27a, M.27b, M.28, Mat28b, Mat28a, Mat29a, Mat29b, O.28a, O.28b, Ost26, P.28, R.26a, R.26b, R.26c, R.26d, R.27b, R.27a, R.28c, R.28a, R.29, Rus28, Rus29, S.26a, S.26b, S.26c, S.26d]. Reviews [S.26e, S.26f, S.27a, S.27b, S.28b, S.28a, S.29c, S.29d, S.29b, S.29a, Sal26, Sch28a, Sch29a, Sch29b, Sch29c, Sch28d, Sch29d, She28a, She28b, Spä26a, Spä26b, Spä26c, Spä28a, Spä28b, Spä28c, Ste29, Sto27, Sto29, Sun26c, Sun27c, Sun27d, Sun27e, Sun28a, Sun28c, TP26a, TP26b, TP28a, TP28b, TP29, T.27, T.28b, T.28c, T.28a, T.29a, T.29b, Tai27, Tho26, W.26, Yon27]. Rhincalanus [Söm29]. Rouge [Le 27a, Le 29]. Rouges [Le 27a]. Rule [Sun29].

Salinity [LC28, Sto28, Pal26]. salt [vR29]. Salzgehalt [Pal26]. same [Joh26a]. Sampler [Lum27, Lum28, Lum29]. Samples [HR27, Lum27]. Sardine [Le 28]. Sars [H.27d]. Sauerstoffanalysen [Buc29b]. Scale [Gra27a]. Scale-Tracings [Gra27a]. Scales [Gra29a, Gra26a]. Schollenbestande [Büc27]. Sea [Lum27, Sho29, Sto28, ZBI28, Hic27, Far28, Hod28, Joh26a, Joh28, Le 27a, Le 29, LA29, Pet26b, TP27, vR29]. Sea-Bottom [ZBI28, Hic27]. Seas [ZBI28]. Seawater [Buc29a]. seine [Ehr27a]. seiner [Ehr27a]. Selected [All26a, All26b]. Selective [BW29, Hod27b, BW27]. Sex [Tes28]. Ships [Lum27]. Sigrid [Ano26f]. Silk [Sei29]. situ [Pet28]. Size [RE26, Hod27a]. Sizes [Dav29]. Skagerak [Mol29]. Slide [Sun29]. Soaps [Atk27]. Some [Hod28, Mol29, Atk26a, Atk26b, BW27, BWH29, Ekm28b]. South [Far27, Har29]. southern [vR29, Hod28, Pet26b, TP27]. Spain [dB27]. Spawning [Hod28, Jen27, Hes27]. Species [Vol28]. specific [Sun26b]. Spring [Joh28]. Stage [Lea29]. Stages [Ruu28]. Staining [Tån27]. Standard [Wim28]. stationärem [Pal26]. Statistics [BWH29, Büc27, Hes27, Hod27a, Sun27a, TP27]. steady [Pal26]. Stickstoff [Buc29b]. Stickstoff- [Buc29b]. Stock [Mol29, TP27, Jen27, Sun27a]. Stocks [Büc27, Hes27, Hod27a, Sun27a, TP27]. Strait [Sch28c, Sch28e]. Straits [dB27]. Stromzustande [Pal26]. Study [ZBI28, Le 27a]. südlichen [vR29]. Summer [Hav28]. Surface [Lum27, Lum28, Lum29, vR29]. Survey [Ekm28b].

Teil [vR29]. Temperatur [Pal26]. Temperature [Hic27, Lum28, Lum29, Lum27, Pal26]. Temperatures [Buc29a, Sun26a]. their [Sch28b]. Theoretical [Ekm28a, Ekm28b]. Thon [Le 27a, Le 29]. Thons [Le 27a]. Thunnus [Le 27a, Le 29]. thynnus [Le 27a, Le 29]. together [Vol28]. Tracings [Gra27a]. Transition [Pet26c, Pet26d]. Transparent [Wim28]. Transplanting [Joh28]. Trawl [BW27, RE26, Ehr27a]. Trawlheringsfang [Ehr27a]. Trawling [Dav29]. Treating [BWH29]. Trouser [RE26]. Tuna [Le 27a, Le 29].

unusual [LA29]. upon [Jen27]. use [Wim28, Buc29b]. using [Atk27].

Valuation [Pet26b]. Variations [Vol28]. Ventilation [Wat29]. Verteilung [Pal26]. Verwendung [Buc29b]. view [Lum28, Tån27]. Volume [Ano26e, Ano27d, Ano28d, Ano29d, Sun26b]. Vorläufige [Wat29]. Vorwiegen [Büc27]. Voyages [Hic28].

Wachstums [Moh27]. Wasserbewegungen [Sch28c, Sch28e]. Wassers [Pal26]. Water [Hic27, Sto28, Atk26a, Atk26b, Pal26, Sch28c, Sch28e]. Waterbottle [Knu29a]. Waters [Har29, Joh28, ME28, Tes28]. which [dB27]. White [ZBI28]. Winter [Joh26a, Gra26a]. Work [dB27, Sch28c]. Works [Pet26b].

year [Hjo26]. years [Jen27]. young [Joh28].

zone [Gra26a]. zur [Sch28c].

References

A:1926:R

[A.26] W. R. G. A. Reviews. ICES Journal of Marine Science, 1(2):161– 162, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

        (electronic). URL http://academic.oup.com/icesjms/article/1/
        2/161/646462.

A:1927:Ra

[A.27a] F. S. A. Reviews. ICES Journal of Marine Science, 2(2):212–213, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

        (electronic). URL http://academic.oup.com/icesjms/article/2/
        2/212/650824.

A:1927:Rb

[A27b] W. R. G. A. Reviews. ICES Journal of Marine Science, 2(3):371–372, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

        (electronic). URL http://academic.oup.com/icesjms/article/2/
        3/371/648924.

Allen:1926:SBMa

[All26a] E. J. Allen. A selected bibliography of marine bionomics and fishery investigation. ICES Journal of Marine Science, 1(1):77–96, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/77/627674.

Allen:1926:SBMb

[All26b] E. J. Allen. A selected bibliography of marine bionomics and fishery investigation. ICES Journal of Marine Science, 1(2):175–193, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/175/646509.

Andersson:1927:R

[And27] K. A. Andersson. Reviews. ICES Journal of Marine Science, 2(2):213– 214, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/213/650826.

Anonymous:1926:A

[Ano26a] Anonymous. ADDENDA. ICES Journal of Marine Science, 1(2): 126, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/126/646450.

Anonymous:1926:CBa

[Ano26b] Anonymous. Current bibliography. ICES Journal of Marine Science, 1(3):293–302, August 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/3/293/632196.

Anonymous:1926:CBb

[Ano26c] Anonymous. Current bibliography. ICES Journal of Marine Science, 1(4):375–383, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/1/4/375/637670.

Anonymous:1926:EN

[Ano26d] Anonymous. Editorial note. ICES Journal of Marine Science, 1 (1):3–4, February 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/1/3/627599.

Anonymous:1926:IV

[Ano26e] Anonymous. Index to volume I. ICES Journal of Marine Science, 1(4): 385–388, November 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/4/385/637674.

Anonymous:1926:MSH

[Ano26f] Anonymous. Miss Sigrid Hansen. ICES Journal of Marine Science, 1(3):196, August 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/3/196/632123.

Anonymous:1926:N

[Ano26g] Anonymous. Notes. ICES Journal of Marine Science, 1(1):75–76, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/75/627673.

Anonymous:1926:NNL

[Ano26h] Anonymous. Notices necrologiques: L’Amiral augusto eduardo neuparth. ICES Journal of Marine Science, 1(3):194–195, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/1/3/
194/632122.

Anonymous:1926:Ra

[Ano26i] Anonymous. Reviews. ICES Journal of Marine Science, 1(1):61–62, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/61/627629.

Anonymous:1926:Rb

[Ano26j] Anonymous. Reviews. ICES Journal of Marine Science, 1(1):62–63, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/62/627633.

Anonymous:1926:Rc

[Ano26k] Anonymous. Reviews. ICES Journal of Marine Science, 1(1):63, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/63/627636.

Anonymous:1926:Rd

[Ano26l] Anonymous. Reviews. ICES Journal of Marine Science, 1(1):64–65, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/64/627641.

Anonymous:1927:CBa

[Ano27a] Anonymous. Current bibliography. ICES Journal of Marine Science, 2(1):101–107, March 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/1/101/633948.

Anonymous:1927:CBb

[Ano27b] Anonymous. Current bibliography. ICES Journal of Marine Science, 2(2):259–267, July 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/2/259/650866.

Anonymous:1927:CBc

[Ano27c] Anonymous. Current bibliography. ICES Journal of Marine Science, 2(3):388–397, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/2/3/388/648978.

Anonymous:1927:IVI

[Ano27d] Anonymous. Index to volume II. ICES Journal of Marine Science, 2(3):398–399, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/2/3/398/648981.

Anonymous:1928:CBa

[Ano28a] Anonymous. Current bibliography. ICES Journal of Marine Science, 3(1):124–131, April 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/1/124/737413.

Anonymous:1928:CBb

[Ano28b] Anonymous. Current bibliography. ICES Journal of Marine Science, 3(2):282–291, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/3/2/282/624212.

Anonymous:1928:CBc

[Ano28c] Anonymous. Current bibliography. ICES Journal of Marine Science, 3(3):412–421, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/3/3/412/690427.

Anonymous:1928:IVI

[Ano28d] Anonymous. Index to volume III. ICES Journal of Marine Science, 3(3):422–424, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/3/3/422/690429.

Anonymous:1928:R

[Ano28e] Anonymous. Reviews. ICES Journal of Marine Science, 3(1):108– 113, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/108/737379.

Anonymous:1929:CBa

[Ano29a] Anonymous. Current bibliography. ICES Journal of Marine Science, 4(1):126–134, March 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/1/126/744035.

Anonymous:1929:CBb

[Ano29b] Anonymous. Current bibliography. ICES Journal of Marine Science, 4(2):252–264, August 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/2/252/631723.

Anonymous:1929:CBc

[Ano29c] Anonymous. Current bibliography. ICES Journal of Marine Science, 4(3):363–371, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/4/3/363/653812.

Anonymous:1929:IVI

[Ano29d] Anonymous. Index to volume IV. ICES Journal of Marine Science, 4(3):372–375, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/4/3/372/653818.

Atkins:1926:QCSa

[Atk26a] W. R. G. Atkins. A quantitative consideration of some factors concerned in plant growth in water. ICES Journal of Marine Science, 1(2):99–126, May 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/2/99/646510.

Atkins:1926:QCSb

[Atk26b] W. R. G. Atkins. A quantitative consideration of some factors concerned in plant growth in water. ICES Journal of Marine Science, 1 (3):197–226, August 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/3/197/632124.

Atkinson:1926:R

[Atk26c] George T. Atkinson. Reviews. ICES Journal of Marine Science, 1(1): 73–74, February 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/1/73/627668.

Atkins:1927:MUC

[Atk27] W. R. G. Atkins. The methods of using copper soaps to preserve fishing nets. ICES Journal of Marine Science, 2(2):144–150, July 1927. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/2/2/144/650803.

B:1926:Ra

[B.26a] J. R. B. Reviews. ICES Journal of Marine Science, 1(3):280–281, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/280/632145.

B:1926:Rb

[B.26b] J. R. B. Reviews. ICES Journal of Marine Science, 1(3):281–282, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/281/632148.

B:1926:Rc

[B.26c] J. R. B. Reviews. ICES Journal of Marine Science, 1(3):282–283, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/282/632151.

B:1926:Rd

[B.26d] J. R. B. Reviews. ICES Journal of Marine Science, 1(4):373–374, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/373/637666.

B:1927:R

[B.27] A. F. B. Reviews. ICES Journal of Marine Science, 2(2):214–216, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/214/650827.

Bckm:1928:R

[Bck28] Bckm. Reviews. ICES Journal of Marine Science, 3(3):393–395, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/393/690394.

BierensDeHaan:1928:R

[Bie28] J. A. Bierens De Haan. Reviews. ICES Journal of Marine Science, 3(3):402–404, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/3/3/402/690410.

Bjerkan:1928:Ra

[Bje28a] Paul Bjerkan. Reviews. ICES Journal of Marine Science, 3(2):269–270, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/269/624187.

Bjerkan:1928:Rb

[Bje28b] Paul Bjerkan. Reviews. ICES Journal of Marine Science, 3(2):270– 272, September 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/2/270/624192.

Blegvad:1927:AOD

[Ble27a] H. Blegvad. An aid to otolith determinations. ICES Journal of Marine Science, 2(2):151–152, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/2/2/151/650805.

Blegvad:1927:Ra

[Ble27b] H. Blegvad. Reviews. ICES Journal of Marine Science, 2(2):220– 222, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/220/650838.

Blegvad:1927:Rb

[Ble27c] H. Blegvad. Reviews. ICES Journal of Marine Science, 2(2):222– 224, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/222/650841.

Blegvad:1928:R

[Ble28] H. Blegvad. Reviews. ICES Journal of Marine Science, 3(3):397–400, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/397/690404.

Blegvad:1929:Ra

[Ble29a] H. Blegvad. Reviews. ICES Journal of Marine Science, 4(1):121–123, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/121/744018.

Blegvad:1929:Rb

[Ble29b] H. Blegvad. Reviews. ICES Journal of Marine Science, 4(1):123–124, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/123/744023.

Blegvad:1929:Rc

[Ble29c] H. Blegvad. Reviews. ICES Journal of Marine Science, 4(1):124–125, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/124/744029.

Buckmann:1927:BSS

[Büc27] Adolf Bückmann. Biological statistics of the stocks of fish: I. Vorwiegen des Jahrgangs 1922 im Schollenbestande der Deutschen Bucht. ICES Journal of Marine Science, 2(1):53–69, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/2/1/53/633958.

Buch:1929:DPS

[Buc29a] K. Buch. On the determination of pH in seawater at different temperatures. ICES Journal of Marine Science, 4(3):267–280, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/267/653744.

Buch:1929:VSS

[Buc29b] Kurt Buch. Die Verwendung von Stickstoff- und Sauerstoffanalysen in der Meeresforschung. (German) [Use of nitrogen and oxygen analysis in marine research]. ICES Journal of Marine Science, 4(2):162–191, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/162/631638.

Buchanan-Wollaston:1925:RCM

[BW25] H. J. Buchanan-Wollaston. A recording current meter. ICES Journal of Marine Science, s1(86):1–14, May 1925. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

Buchanan-Wollaston:1927:SAT

[BW27] H. J. Buchanan-Wollaston. On the selective action of a trawl net, with some remarks on selective action of drift nets. ICES Journal of Marine Science, 2(3):343–355, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/2/3/343/648911.

Buchanan-Wollaston:1929:SAG

[BW29] H. J. Buchanan-Wollaston. The selective action of the gelder cod-end and that of other cod-ends compared. ICES Journal of Marine Science, 4(3):300–308, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/4/3/300/653754.

Buchanan-Wollaston:1929:NMT

[BWH29] H. J. Buchanan-Wollaston and William C. Hodgson. A new method of treating frequency curves in fishery statistics, with some results. ICES Journal of Marine Science, 4(2):207–225, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/2/207/631653.

C:1926:Ra

[C.26a] J. N. C. Reviews. ICES Journal of Marine Science, 1(3):286–289, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/286/632166.

C:1926:Rb

[C.26b] J. N. C. Reviews. ICES Journal of Marine Science, 1(4):354–356, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/354/637597.

C:1926:Rc

[C.26c] J. N. C. Reviews. ICES Journal of Marine Science, 1(4):356–357, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/356/637602.

C:1928:R

[C.28] R. S. C. Reviews. ICES Journal of Marine Science, 3(2):257–261, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/257/624170.

C:1929:R

[C.29] J. N. C. Reviews. ICES Journal of Marine Science, 4(3):344–345, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/344/653775.

Carruthers:1926:NCM

[Car26] M. Sc. J. N. Carruthers. A new current measuring instrument for the purposes of fishery research. ICES Journal of Marine Science, 1 (2):127–139, May 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/2/127/646451.

Carruthers:1928:NDB

[Car28] D. Sc. J. N. Carruthers. New drift bottles for the investigation of currents in connection with fishery research. ICES Journal of Marine Science, 3(2):194–205, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/3/2/194/624147.

Carruthers:1929:R

[Car29] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 4(2): 230–234, August 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/2/230/631661.

Davis:1926:R

[Dav26] F. M. Davis. Reviews. ICES Journal of Marine Science, 1(1):69–70, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/69/627660.

Davis:1929:PNE

[Dav29] F. M. Davis. Preliminary note on experimental trawling with codend meshes of different sizes. ICES Journal of Marine Science, 4(3): 287–299, December 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/3/287/653752.

deBuen:1927:AWW

[dB27] Rafael de Buen. An account of the work which Spain has carried out in the Atlantic Ocean and the Straits of Gibraltar. ICES Journal of Marine Science, 2(3):269–308, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/2/3/269/648898.

Dannevig:1929:RLL

[DM29] Alf Dannevig and A. Meek. The rearing of lobster larvae at Flødevigen. ICES Journal of Marine Science, 4(1):105–106, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/1/105/743929.

E:1929:R

[E.29] T. E. Reviews. ICES Journal of Marine Science, 4(1):118–119, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/118/744006.

Edser:1926:R

[Eds26] T. Edser. Reviews. ICES Journal of Marine Science, 1(3):291–292, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/291/632193.

Eh:1926:R

[Eh.26] Eh. Reviews. ICES Journal of Marine Science, 1(4):371–373, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/371/637662.

Eh:1927:R

[Eh.27] Eh. Reviews. ICES Journal of Marine Science, 2(2):230–231, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/2/2/
230/650854.

Ehrenbaum:1927:TSE

[Ehr27a] E. Ehrenbaum. Der Trawlheringsfang, seine Entwicklung, seine Bedeutung und die Erforschung seiner Grundlagen. (German) [The trawl herring catch, its development, its importance and the research into its basics]. ICES Journal of Marine Science, 2(2):113–131, July 1927. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/2/2/113/650799.

Ehrenbaum:1927:R

[Ehr27b] E. Ehrenbaum. Reviews. ICES Journal of Marine Science, 2(3): 382–384, December 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/3/382/648967.

Ekman:1928:NTC

[Ekm28a] V. Walfrid Ekman. Note on the theoretical courses of ocean-currents. ICES Journal of Marine Science, 3(1):104–107, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/3/1/104/737376.

Ekman:1928:SST

[Ekm28b] V. Walfrid Ekman. A survey of some theoretical investigations on ocean-currents. ICES Journal of Marine Science, 3(3):295–327, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/295/690340.

F:1926:Rd

[F.26a] E. F. Reviews. ICES Journal of Marine Science, 1(4):363–365, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/363/637637.

F:1926:Re

[F.26b] E. F. Reviews. ICES Journal of Marine Science, 1(4):365–366, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/365/637640.

F:1926:Ra

[F.26c] G. P. F. Reviews. ICES Journal of Marine Science, 1(3):283–284, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/283/632156.

F:1926:Rb

[F.26d] G. P. F. Reviews. ICES Journal of Marine Science, 1(4):359, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/359/637612.

F:1926:Rc

[F.26e] G. P. F. Reviews. ICES Journal of Marine Science, 1(4):360, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/360/637617.

F:1927:Ra

[F.27a] G. P. F. Reviews. ICES Journal of Marine Science, 2(1):82–83, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/82/633970.

F:1927:Rb

[F.27b] G. P. F. Reviews. ICES Journal of Marine Science, 2(1):83–85, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/83/633973.

F:1927:Rc

[F.27c] G. P. F. Reviews. ICES Journal of Marine Science, 2(1):85, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/85/633974.

F:1927:Rd

[F.27d] G. P. F. Reviews. ICES Journal of Marine Science, 2(3):378–379, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/378/648945.

F:1927:Re

[F.27e] G. P. F. Reviews. ICES Journal of Marine Science, 2(3):379–380, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/379/648952.

F:1928:Rb

[F.28a] E. F. Reviews. ICES Journal of Marine Science, 3(3):391–393, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/391/690391.

F:1928:Ra

[F.28b] G. P. F. Reviews. ICES Journal of Marine Science, 3(3):389–390, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/389/690383.

F:1928:Rc

[F.28c] G. P. F. Reviews. ICES Journal of Marine Science, 3(3):400–402, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/400/690407.

F:1929:Ra

[F.29a] E. F. Reviews. ICES Journal of Marine Science, 4(1):119–120, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/119/744013.

F:1929:Rd

[F.29b] E. F. Reviews. ICES Journal of Marine Science, 4(3):350–352, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/350/653790.

F:1929:Rb

[F.29c] G. P. F. Reviews. ICES Journal of Marine Science, 4(2):243–244, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/243/631709.

F:1929:Rc

[F.29d] G. P. F. Reviews. ICES Journal of Marine Science, 4(2):244–245, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/244/631713.

Farran:1927:RCF

[Far27] G. P. Farran. The reproduction of Calanus finmarchicus off the South Coast of Ireland. ICES Journal of Marine Science, 2(2):132–143, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/132/650800.

Farran:1928:NGR

[Far28] G. P. Farran. Note on the growth-rate of herrings in the Irish Sea. ICES Journal of Marine Science, 3(3):360–366, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/3/3/360/690347.

Gilson:1926:R

[Gil26] G. Gilson. Reviews. ICES Journal of Marine Science, 1(1):55–59, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/55/627622.

Gilson:1927:R

[Gil27] Gilson. Reviews. ICES Journal of Marine Science, 2(1):93–96, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/93/633991.

Gilson:1928:R

[Gil28] G. Gilson. Reviews. ICES Journal of Marine Science, 3(2):248–251, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/248/624158.

Graham:1926:PMD

[Gra26a] M. A. Michael Graham. A precise method for determining the first “winter” zone in cod scales. ICES Journal of Marine Science, 1(4): 344–352, November 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/4/344/637590.

Graham:1926:Ra

[Gra26b] Michael Graham. Reviews. ICES Journal of Marine Science, 1 (2):160–161, May 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/2/160/646459.

Graham:1926:Rb

[Gra26c] Michael Graham. Reviews. ICES Journal of Marine Science, 1(4): 361–363, November 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/4/361/637631.

Graham:1927:MRG

[Gra27a] Michael Graham. A machine for reproducing gadoid scale-tracings. ICES Journal of Marine Science, 2(2):153–157, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/2/2/153/650807.

Graham:1927:Ra

[Gra27b] Michael Graham. Reviews. ICES Journal of Marine Science, 2(1): 89, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/89/633984.

Graham:1927:Rb

[Gra27c] Michael Graham. Reviews. ICES Journal of Marine Science, 2(2): 232, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/232/650860.

Gran:1928:R

[Gra28] H. H. Gran. Reviews. ICES Journal of Marine Science, 3(2):251–252, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/251/624160.

Graham:1929:CPF

[Gra29a] Michael Graham. On the collection and preparation of fish scales. ICES Journal of Marine Science, 4(3):339–340, December 1929. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/4/3/339/653761.

Graham:1929:R

[Gra29b] Michael Graham. Reviews. ICES Journal of Marine Science, 4(3): 359–362, December 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/3/359/653809.

H:1926:R

[H.26] W. C. H. Reviews. ICES Journal of Marine Science, 1(4):370–371, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/370/637658.

H:1927:Rd

[H.27a] C. F. H. Reviews. ICES Journal of Marine Science, 2(3):380–381, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/380/648956.

H:1927:Rb

[H.27b] H. W. H. Reviews. ICES Journal of Marine Science, 2(3):372–374, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/372/648929.

H:1927:Rc

[H.27c] H. W. H. Reviews. ICES Journal of Marine Science, 2(3):374–376, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/374/648934.

H:1927:GOS

[H.27d] J. H. Georg Ossian Sars. (1837–1927). ICES Journal of Marine Science, 2(2):113, July 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/2/113/650798.

H:1927:Ra

[H.27e] W. C. H. Reviews. ICES Journal of Marine Science, 2(2):224–226, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/224/650845.

H:1928:Rb

[H.28a] C. F. H. Reviews. ICES Journal of Marine Science, 3(3):404–405, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/404/690416.

H:1928:Ra

[H28b] H. W. H. Reviews. ICES Journal of Marine Science, 3(3):388–389, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/388/690380.

H:1929:Rc

[H.29a] E. H. Reviews. ICES Journal of Marine Science, 4(3):347–349, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/347/653783.

H:1929:Rd

[H.29b] E. H. Reviews. ICES Journal of Marine Science, 4(3):349–350, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/349/653786.

H:1929:Rb

[H.29c] G. W. H. Reviews. ICES Journal of Marine Science, 4(1):116–118, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/116/743999.

H:1929:Ra

[H.29d] H. W. H. Reviews. ICES Journal of Marine Science, 4(1):109–110, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/109/743945.

H:1929:Re

[H.29e] W. C. H. Reviews. ICES Journal of Marine Science, 4(3):352–353, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/352/653793.

Harris:1927:R

[Har27] G. W. Harris. Reviews. ICES Journal of Marine Science, 2(1):88–89, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/88/633980.

Harvey:1929:HWS

[Har29] H. W. Harvey. Hydrodynamics of the waters South East of Ireland. ICES Journal of Marine Science, 4(1):80–92, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/1/80/744052.

Havinga:1928:DRG

[Hav28] B. Havinga. The daily rate of growth of oysters during summer. ICES Journal of Marine Science, 3(2):231–245, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/3/2/231/624153.

Heldt:1926:R

[Hel26] H. Heldt. Reviews. ICES Journal of Marine Science, 1(1):68–69, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/68/627653.

Hessle:1927:BSS

[Hes27] Chr. Hessle. Biological statistics of the stocks of fish: II. Biological statistics regarding the Baltic autumn-spawning herring. ICES Journal of Marine Science, 2(2):169–172, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/2/2/169/650814.

Hickling:1927:IRA

[Hic27] C. F. Hickling. An improvised reversing apparatus for finding the temperature of the water at the sea-bottom. ICES Journal of Marine Science, 2(2):158–160, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/2/2/158/650808.

Hickling:1928:FEV

[Hic28] C. F. Hickling. The fleetwood exploratory voyages for hake. ICES Journal of Marine Science, 3(1):70–89, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/3/1/70/737419.

Hjort:1926:FYC

[Hjo26] Johan Hjort. Fluctuations in the year classes of important food fishes. ICES Journal of Marine Science, 1(1):5–38, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/1/1/5/627619.

Hodgson:1926:R

[Hod26] W. C. Hodgson. Reviews. ICES Journal of Marine Science, 1(2):163– 167, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/163/646482.

Hodgson:1927:BSS

[Hod27a] B. Sc. William C. Hodgson. Biological statistics of the stocks of fish: II. Size and age composition of the East Anglian autumn herring. ICES Journal of Marine Science, 2(1):69–74, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/2/1/69/633960.

Hodgson:1927:PNE

[Hod27b] W. C. Hodgson. Preliminary note on experiments concerning the selective action of drift nets. ICES Journal of Marine Science, 2(3): 356–360, December 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/3/356/648914.

Hodgson:1928:SSP

[Hod28] M. Sc. W. C. Hodgson. Some spawning places of the herring of the Southern North Sea. ICES Journal of Marine Science, 3(2):224–230, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/224/624151.

Hjort:1927:MAC

[HR27] Johan Hjort and Johan T. Ruud. A method for the analysis and comparison of plankton samples. ICES Journal of Marine Science, 2(1):28– 37, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/28/633949.

Hunt:1927:R

[Hun27] O. D. Hunt. Reviews. ICES Journal of Marine Science, 2(3):381–382, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/381/648960.

J:1926:FCH

[J.26a] J. J. Frequency curves and herring investigation. ICES Journal of Marine Science, 1(3):276–277, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/1/3/276/632135.

J:1926:R

[J.26b] J. J. Reviews. ICES Journal of Marine Science, 1(3):278, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/278/632139.

J:1927:R

[J.27] J. T. J. Reviews. ICES Journal of Marine Science, 2(3):384–385, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/384/648970.

J:1929:Ra

[J.29a] J. T. J. Reviews. ICES Journal of Marine Science, 4(3):356–358, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/356/653803.

J:1929:Rb

[J.29b] J. T. J. Reviews. ICES Journal of Marine Science, 4(3):358–359, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/358/653807.

Jacobsen:1928:R

[Jac28] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 3(3):382– 384, December 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/3/382/690372.

Jensen:1927:IQS

[Jen27] Aage J. C. Jensen. On the influence of the quantity of spawning herrings upon the stock of the following years. ICES Journal of Marine Science, 2(1):44–49, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/2/1/44/633954.

Jespersen:1926:ADE

[Jes26] P. Jespersen. Age determination of eels from Norfolk and Cumberland. ICES Journal of Marine Science, 1(3):227–234, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/1/3/227/632128.

Jespersen:1928:Ra

[Jes28a] P. Jespersen. Reviews. ICES Journal of Marine Science, 3(1):116– 117, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/116/737393.

Jespersen:1928:Rb

[Jes28b] P. Jespersen. Reviews. ICES Journal of Marine Science, 3(3):390–391, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/390/690387.

Johansen:1926:RQH

[Joh26a] A. C. Johansen. On the remarkable quantities of haddock in the Belt Sea during the winter of 1925–26, and causes leading to the same. ICES Journal of Marine Science, 1(2):140–156, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/1/2/140/646453.

Johansen:1926:R

[Joh26b] A. C. Johansen. Reviews. ICES Journal of Marine Science, 1(2):167– 169, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/167/646486.

Johansen:1927:MH

[Joh27] A. C. Johansen. On the migrations of the herring. ICES Journal of Marine Science, 2(1):3–27, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/2/1/3/633951.

Johansen:1928:ETY

[Joh28] A. C. Johansen. Experiments in transplanting young plaice from the North Sea and the Limfjord to the Belt Sea and adjacent waters in the spring of 1928. ICES Journal of Marine Science, 3(3):367–370, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/367/690350.

Knudsen:1929:FRW

[Knu29a] Martin Knudsen. A frameless reversing waterbottle. ICES Journal of Marine Science, 4(2):192–193, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/4/2/192/631643.

Knudsen:1929:Ra

[Knu29b] Martin Knudsen. Reviews. ICES Journal of Marine Science, 4(1): 108–109, March 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/1/108/743938.

Knudsen:1929:Rb

[Knu29c] Martin Knudsen. Reviews. ICES Journal of Marine Science, 4(3):343, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/343/653767.

Kunne:1929:R

[Kün29] Künne. Reviews. ICES Journal of Marine Science, 4(3):353–354, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/353/653797.

L:1926:Ra

[L.26a] J. R. L. Reviews. ICES Journal of Marine Science, 1(1):59–61, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/59/627625.

L:1926:Rb

[L.26b] J. R. L. Reviews. ICES Journal of Marine Science, 1(2):171–173, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/171/646502.

L:1927:Ra

[L.27a] J. R. L. Reviews. ICES Journal of Marine Science, 2(1):75–78, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/75/633962.

L:1927:Rb

[L.27b] J. R. L. Reviews. ICES Journal of Marine Science, 2(2):209–212, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/209/650821.

L:1928:R

[L.28] J. R. L. Reviews. ICES Journal of Marine Science, 3(3):385–388, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/385/690377.

L:1929:Ra

[L.29a] J. R. L. Reviews. ICES Journal of Marine Science, 4(2):236–238, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/236/631668.

L:1929:Rb

[L.29b] M. V. L. Reviews. ICES Journal of Marine Science, 4(2):242–243, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/242/631705.

L:1929:Rc

[L.29c] M. V. L. Reviews. ICES Journal of Marine Science, 4(3):346–347, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/346/653780.

Lumby:1929:UMA

[LA29] J. R. Lumby and G. T. Atkinson. On the unusual mortality amongst fish during March and April 1929, in the North Sea. ICES Journal of Marine Science, 4(3):309–332, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/3/309/653755.

Lumby:1928:EDI

[LC28] J. R. Lumby and D. Sc. J. N. Carruthers. An experiment dealing with the interpretation of salinity distribution charts. ICES Journal of Marine Science, 3(2):176–190, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/3/2/176/624137.

LeGall:1927:CEB

[Le 27a] J. Le Gall. Contribution à l’étude de la biologie du thon rouge (Thunnus thynnus L.): Sur la présence de thons rouges en Mer du nord et dans l’Atlantique nord-est. (French) [Contribution to the study of the biology of bluefin tuna (Thunnus thynnus L.): On the presence of bluefin tuna in the North Sea and in the North-East Atlantic]. ICES Journal of Marine Science, 2(3):309–329, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/2/3/309/648903.

LeGall:1927:C

[Le 27b] Jean Le Gall. Correspondence. ICES Journal of Marine Science, 2 (2):233–235, July 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/2/233/650861.

LeGall:1928:NOB

[Le 28] Jean Le Gall. Notes et observations sur la biologie de la sardine de la Manche. (French) [notes and observations on the biology of the English Channel sardine]. ICES Journal of Marine Science, 3(2):206–223, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/206/624149.

LeGall:1929:TRT

[Le 29] J. Le Gall. Le thon rouge (Thunnus thynnus L.) en Mer du Nord et dans l’Atlantique nord est. (French) [The bluefin tuna (Thunnus thynnus L.) in the North Sea and in the Atlantic North East]. ICES Journal of Marine Science, 4(2):200–206, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/2/200/631648.

Lea:1929:OSL

[Lea29] Einar Lea. The oceanic stage in the life history of the Norwegian herring. ICES Journal of Marine Science, 4(1):3–42, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/4/1/3/744037.

Lumby:1927:SSA

[Lum27] J. R. Lumby. The surface sampler, an apparatus for the collection of samples from the sea surface from ships in motion: With a note on surface temperature observations. ICES Journal of Marine Science, 2(3): 332–342, December 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/3/332/648908.

Lumby:1928:MSS

[Lum28] J. R. Lumby. Modification of the surface sampler with a view to the improvement of temperature observation. ICES Journal of Marine Science, 3(3):340–350, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/3/3/340/690343.

Lumby:1929:SST

[Lum29] J. R. Lumby. A surface sampler for temperature observations. ICES Journal of Marine Science, 4(3):281–283, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/3/281/653748.

M:1926:R

[M.26] D. J. M. Reviews. ICES Journal of Marine Science, 1(3):284–286, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/284/632160.

M:1927:Ra

[M.27a] M. Reviews. ICES Journal of Marine Science, 2(1):96–99, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/2/1/96/633994.

M:1927:Rb

[M.27b] D. J. M. Reviews. ICES Journal of Marine Science, 2(3):361–366, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/361/648917.

M:1928:R

[M.28] A. M. Reviews. ICES Journal of Marine Science, 3(3):404, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/404/690414.

M:1929:FH

[M.29] W. M. Friedrich Heincke (1852–1929). ICES Journal of Marine Science, 4(2):145, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/4/2/145/631632.

Matthews:1928:Rb

[Mat28a] D. J. Matthews. Reviews. ICES Journal of Marine Science, 3(3): 384–385, December 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/3/384/690374.

Matthews:1928:Ra

[Mat28b] Donald J. Matthews. Reviews. ICES Journal of Marine Science, 3(2): 278–279, September 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/2/278/624210.

Matthews:1929:Ra

[Mat29a] D. J. Matthews. Reviews. ICES Journal of Marine Science, 4(1): 107–108, March 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/1/107/743933.

Matthews:1929:Rb

[Mat29b] D. J. Matthews. Reviews. ICES Journal of Marine Science, 4(2):238, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/238/631679.

Matthews:1928:RMC

[ME28] Donald J. Matthews and B. A. Ellis. The ratio of magnesium to chlorine in the waters of the Eastern Mediterranean and Gulf of Aden. ICES Journal of Marine Science, 3(2):191–193, September 1928. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/3/2/191/624145.

Mohr:1927:BAW

[Moh27] Erna W. Mohr. Bibliographie der Alters- und Wachstums Bestimmung bei Fischen. (German) [Bibliography of age and growth determination in fish]. ICES Journal of Marine Science, 2(2):236–258, July 1927. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/2/2/236/650864.

Molander:1929:SOC

[Mol29] Arvid R. Molander. Some observations concerning the stock of haddock in the Skagerak and Kattegat. ICES Journal of Marine Science, 4(3): 333–338, December 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/3/333/653759.

O:1928:Ra

[O.28a] A. P. O. Reviews. ICES Journal of Marine Science, 3(1):115–116, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/115/737389.

O:1928:Rb

[O.28b] A. P. O. Reviews. ICES Journal of Marine Science, 3(1):116, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/116/737391.

Orton:1928:C

[Ort28] J. H. Orton. Correspondence. ICES Journal of Marine Science, 3 (3):410, December 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/3/410/690423.

Ostenfeld:1926:R

[Ost26] C. H. Ostenfeld. Reviews. ICES Journal of Marine Science, 1(4): 357–358, November 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/4/357/637607.

P:1928:R

[P.28] B. W. P. Reviews. ICES Journal of Marine Science, 3(1):119–121, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/119/737409.

Palmen:1926:VTS

[Pal26] E. Palmén. Über die Verteilung von Temperatur und Salzgehalt des Wassers im Finnischen Meerbusen bei stationärem Stromzustande. (German) [On the distribution of temperature and salinity of the water in the Gulf of Finland with steady flow conditions]. ICES Journal of Marine Science, 1(4):327–343, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/1/4/327/637587.

Petersen:1926:C

[Pet26a] C. G. Joh. Petersen. Correspondence. ICES Journal of Marine Science, 1(4):353, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/1/4/353/637594.

Petersen:1926:WCV

[Pet26b] C. G. Joh. Petersen. On works concerning the valuation of the Southern North Sea. ICES Journal of Marine Science, 1(2):157–160, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/157/646457.

Pettersson:1926:CFM

[Pet26c] Otto Pettersson. Currents and fish-migrations in the transition area. ICES Journal of Marine Science, 1(4):322–326, November 1926. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/1/4/322/637584.

Pettersson:1926:HCF

[Pet26d] Otto Pettersson. Hydrography, climate and fisheries in the transition area. ICES Journal of Marine Science, 1(4):305–321, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/305/637582.

Pettersson:1928:AQM

[Pet28] Vilhelm Irgens Pettersson. Apparatus for quantitative measurements of plankton in situ. ICES Journal of Marine Science, 3(3):351–359, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/351/690346.

R:1926:Ra

[R.26a] E. S. R. Reviews. ICES Journal of Marine Science, 1(1):65–66, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/65/627644.

R:1926:Rb

[R.26b] E. S. R. Reviews. ICES Journal of Marine Science, 1(1):66–67, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/66/627646.

R:1926:Rc

[R.26c] E. S. R. Reviews. ICES Journal of Marine Science, 1(2):171, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/171/646497.

R:1926:Rd

[R.26d] F. S. R. Reviews. ICES Journal of Marine Science, 1(3):279–280, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/279/632143.

R:1927:Rb

[R.27a] F. S. R. Reviews. ICES Journal of Marine Science, 2(2):217–220, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/217/650835.

R:1927:Ra

[R.27b] H. C. R. Reviews. ICES Journal of Marine Science, 2(1):80–82, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/80/633968.

R:1928:Rb

[R.28a] C. H. R. Reviews. ICES Journal of Marine Science, 3(2):277–278, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/277/624206.

R:1928:CGJ

[R.28b] E. S. R. C. G. Joh. Petersen. (1860–1928). ICES Journal of Marine Science, 3(2):139, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/3/2/139/624129.

R:1928:Ra

[R.28c] H. C. R. Reviews. ICES Journal of Marine Science, 3(2):274–277, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/274/624204.

R:1929:R

[R.29] H. C. R. Reviews. ICES Journal of Marine Science, 4(1):110–112, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/110/743950.

Russell:1926:RBC

[RE26] E. S. Russell and T. Edser. The relation between cod-end mesh and size of fish caught: Preliminary experiments with the trouser trawl. ICES Journal of Marine Science, 1(1):39–54, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/1/1/39/627608.

Roberts:1926:OP

[Rob26] B. A. C. H. Roberts. Oil pollution. ICES Journal of Marine Science, 1 (3):245–275, August 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/3/245/632134.

Russell:1928:R

[Rus28] F. S. Russell. Reviews. ICES Journal of Marine Science, 3(2):254–255, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/254/624165.

Russell:1929:R

[Rus29] F. S. Russell. Reviews. ICES Journal of Marine Science, 4(3):345–346, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/345/653778.

Ruud:1928:FLS

[Ruu28] Johan T. Ruud. Fluctuations in the larval stages of Euphausiidae off Möre, 1926–1927. ICES Journal of Marine Science, 3(1):98–101, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/98/737427.

S:1926:Ra

[S.26a] R. S. Reviews. ICES Journal of Marine Science, 1(2):162, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/1/2/
162/646469.

S:1926:Rb

[S.26b] R. S. Reviews. ICES Journal of Marine Science, 1(2):162–163, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/162/646473.

S:1926:Rc

[S.26c] R. S. Reviews. ICES Journal of Marine Science, 1(2):163, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/1/2/
163/646477.

S:1926:Rd

[S.26d] R. S. Reviews. ICES Journal of Marine Science, 1(3):289, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/289/632172.

S:1926:Re

[S.26e] R. S. Reviews. ICES Journal of Marine Science, 1(3):289–290, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/289/632176.

S:1926:Rf

[S.26f] R. S. Reviews. ICES Journal of Marine Science, 1(3):290, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/290/632183.

S:1927:Ra

[S.27a] R. E. S. Reviews. ICES Journal of Marine Science, 2(2):216–217, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/216/650831.

S:1927:Rb

[S.27b] R. E. S. Reviews. ICES Journal of Marine Science, 2(3):376–378, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/376/648941.

S:1928:Rb

[S.28a] H. P. S. Reviews. ICES Journal of Marine Science, 3(1):121–123, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/121/737411.

S:1928:Ra

[S.28b] O. S. Reviews. ICES Journal of Marine Science, 3(1):115, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/3/1/
115/737386.

S:1929:Rd

[S.29a] H. P. S. Reviews. ICES Journal of Marine Science, 4(2):249–251, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/249/631721.

S:1929:Rc

[S.29b] M. S. Reviews. ICES Journal of Marine Science, 4(2):248–249, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/248/631719.

S:1929:Ra

[S.29c] O. S. Reviews. ICES Journal of Marine Science, 4(1):116, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/4/1/
116/743992.

S:1929:Rb

[S.29d] R. E. S. Reviews. ICES Journal of Marine Science, 4(2):241–242, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/241/631700.

Salisbury:1926:R

[Sal26] E. J. Salisbury. Reviews. ICES Journal of Marine Science, 1(4):360– 361, November 1926. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/1/4/360/637625.

Schmidt:1927:ELF

[Sch27a] Johs. Schmidt. Eel larvæ in the Faroe Channel. ICES Journal of Marine Science, 2(1):38–43, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/2/1/38/633953.

Schott:1927:CN

[Sch27b] G. Schott. Current notes. ICES Journal of Marine Science, 2(3): 386–387, December 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/3/386/648975.

Schnakenbeck:1928:R

[Sch28a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 3(2):265– 269, September 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/2/265/624182.

Schnakenbeck:1928:OIB

[Sch28b] W. Schnakenbeck. Otolithen und ihre Beurteilung. (German) [Otoliths and their assessment]. ICES Journal of Marine Science, 3(1):90–93, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/90/737422.

Schott:1928:NAW

[Sch28c] G. Schott. Notiz zur Arbeit: Die Wasserbewegungen im Gebiete der Gibraltarstrasse. (German) [Note on work: Water movements in the area of the Gibraltar Strait]. ICES Journal of Marine Science, 3(3): 411, December 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/3/411/690424. See [Sch28e].

Schott:1928:R

[Sch28d] G. Schott. Reviews. ICES Journal of Marine Science, 3(2):246–248, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/246/624155.

Schott:1928:WIG

[Sch28e] Gerhard Schott. Die Wasserbewegungen im Gebiete der Gibraltarstrasse. (German) [Water movements in the area of the Gibraltar Strait]. ICES Journal of Marine Science, 3(2):139–175, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/

2/139/624132. See comment [Sch28c].

Schnakenbeck:1929:Ra

[Sch29a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 4(1): 112–114, March 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/1/112/743959.

Schnakenbeck:1929:Rb

[Sch29b] Schnakenbeck. Reviews. ICES Journal of Marine Science, 4(1): 114–116, March 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/1/114/743986.

Schnakenbeck:1929:Rc

[Sch29c] Schnakenbeck. Reviews. ICES Journal of Marine Science, 4(2): 247–248, August 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/2/247/631717.

Schulz:1929:R

[Sch29d] Bruno Schulz. Reviews. ICES Journal of Marine Science, 4(2): 234–236, August 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/2/234/631664.

Seiwell:1929:PCS

[Sei29] H. R. Seiwell. Patterns for conical silk plankton-nets of one meter and half-meter diameters. ICES Journal of Marine Science, 4(1):99–103, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/99/744064.

Sherwood:1928:Ra

[She28a] H. P. Sherwood. Reviews. ICES Journal of Marine Science, 3(3): 405–407, December 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/3/405/690419.

Sherwood:1928:Rb

[She28b] H. P. Sherwood. Reviews. ICES Journal of Marine Science, 3(3): 407–409, December 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/3/407/690420.

Shoulejkin:1929:MOS

[Sho29] Was. W. Shoulejkin. A maregraph for the open sea. ICES Journal of Marine Science, 4(2):194–199, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/4/2/194/631645.

Somme:1929:NNL

[Söm29] Jacob D. Sömme. Note on the northern limit of the distribution of Rhincalanus nasutus Giesbrecht. ICES Journal of Marine Science, 4(3):284–286, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/4/3/284/653751.

Sparck:1926:Ra

[Spä26a] R. Spärck. Reviews. ICES Journal of Marine Science, 1(4):367, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/367/637645.

Sparck:1926:Rb

[Spä26b] R. Spärck. Reviews. ICES Journal of Marine Science, 1(4):367–368, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/367/637649.

Sparck:1926:Rc

[Spä26c] R. Spärck. Reviews. ICES Journal of Marine Science, 1(4):368–370, November 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
4/368/637654.

Sparck:1928:Ra

[Spä28a] R. Spärck. Reviews. ICES Journal of Marine Science, 3(2):272, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/272/624195.

Sparck:1928:Rb

[Spä28b] R. Spärck. Reviews. ICES Journal of Marine Science, 3(2):273, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/273/624198.

Sparck:1928:Rc

[Spä28c] R. Spärck. Reviews. ICES Journal of Marine Science, 3(2):273, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/273/624201.

Stahlberg:1927:DAE

[Sta27] Walter Stahlberg. Die Deutsche Atlantische Expedition. (German) [The German Atlantic Expedition]. ICES Journal of Marine Science, 2(2):184–208, July 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/2/184/650819.

Stephensen:1929:R

[Ste29] K. Stephensen. Reviews. ICES Journal of Marine Science, 4(2): 238–240, August 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/2/238/631688.

Storrow:1927:R

[Sto27] B. Storrow. Reviews. ICES Journal of Marine Science, 2(1):85–87, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/85/633977.

StottBAFInstP:1928:ADS

[Sto28] Verney Stott, B.A., F.Inst.P. An apparatus for determining the salinity of sea water. ICES Journal of Marine Science, 3(3):328–339, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/328/690342.

Stott:1929:R

[Sto29] Verney Stott. Reviews. ICES Journal of Marine Science, 4(3):344, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/344/653770.

Sund:1926:CT

[Sun26a] Oscar Sund. Correction of temperatures. ICES Journal of Marine Science, 1(3):242–244, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/1/3/242/632133.

Sund:1926:GCS

[Sun26b] Oscar Sund. Graphical calculation of specific volume and dynamic depth. ICES Journal of Marine Science, 1(3):235–241, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/235/632131.

Sund:1926:R

[Sun26c] Oscar Sund. Reviews. ICES Journal of Marine Science, 1(3):290–291, August 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
3/290/632188.

Sund:1927:BSS

[Sun27a] Oscar Sund. Biological statistics of the stocks of fish: I. The Arcto-Norwegian cod stock. ICES Journal of Marine Science, 2(2):161–169, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/161/650811.

Sund:1927:C

[Sun27b] Oscar Sund. Correspondence. ICES Journal of Marine Science, 2 (1):100, March 1927. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/2/1/100/633943.

Sund:1927:Ra

[Sun27c] Oscar Sund. Reviews. ICES Journal of Marine Science, 2(1):78–79, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/78/633964.

Sund:1927:Rb

[Sun27d] Oscar Sund. Reviews. ICES Journal of Marine Science, 2(1):79–80, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/79/633965.

Sund:1927:Rc

[Sun27e] Oscar Sund. Reviews. ICES Journal of Marine Science, 2(1):89–91, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/89/633987.

Sund:1928:Ra

[Sun28a] O. Sund. Reviews. ICES Journal of Marine Science, 3(1):113–115, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/113/737383.

Sund:1928:NCC

[Sun28b] Oscar Sund. A new colour comparator. ICES Journal of Marine Science, 3(1):102–103, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/3/1/102/737372.

Sund:1928:Rb

[Sun28c] Oscar Sund. Reviews. ICES Journal of Marine Science, 3(2):252–254, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/252/624162.

Sund:1929:OSR

[Sun29] Oscar Sund. An oceanographical slide rule: a new apparatus for calculating oceanographical data. ICES Journal of Marine Science, 4(1):93– 98, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
1/93/744056.

T:1927:R

[T.27] J. J. T. Reviews. ICES Journal of Marine Science, 2(2):226–230, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
2/226/650850.

T:1928:Rc

[T.28a] H. T. Reviews. ICES Journal of Marine Science, 3(2):261–263, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/261/624173.

T:1928:Ra

[T.28b] J. J. T. Reviews. ICES Journal of Marine Science, 3(1):117–119, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/117/737407.

T:1928:Rb

[T.28c] J. J. T. Reviews. ICES Journal of Marine Science, 3(2):255–257, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/255/624167.

T:1929:Ra

[T.29a] H. T. Reviews. ICES Journal of Marine Science, 4(2):245–247, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
2/245/631716.

T:1929:Rb

[T.29b] J. J. T. Reviews. ICES Journal of Marine Science, 4(3):341–343, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/341/653763.

Tait:1927:R

[Tai27] J. B. Tait. Reviews. ICES Journal of Marine Science, 2(3):366–371, December 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
3/366/648920.

Taaning:1927:DSF

[Tån27] Å. Vedel Tåning. Directions for staining fish with a view to racial investigations. ICES Journal of Marine Science, 2(1):50–52, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/2/1/50/633957.

Thompson:1929:C

[TE29] W. F. Thompson and T. E. Correspondence. ICES Journal of Marine Science, 4(2):226–229, August 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/4/2/226/631655.

Tesch:1928:SGI

[Tes28] J. J. Tesch. On sex and growth investigations on the freshwater eel in Dutch waters. ICES Journal of Marine Science, 3(1):52–69, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
1/52/737417.

Thompson:1926:R

[Tho26] H. Thompson. Reviews. ICES Journal of Marine Science, 1(2):169– 171, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/169/646491.

Thompson:1929:AMF

[Tho29a] D’Arcy W. Thompson. Apparatus for measuring fish. ICES Journal of Marine Science, 4(1):104, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/4/1/104/743924.

Thompson:1929:RHF

[Tho29b] William F. Thompson. The regulation of the halibut fisheries of the Pacific Coast of North America. ICES Journal of Marine Science, 4 (2):145–161, August 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/2/145/631635.

T-P:1926:Ra

[TP26a] D. E. T-P. Reviews. ICES Journal of Marine Science, 1(1):70–72, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/70/627665.

T-P:1926:Rb

[TP26b] D. E. T-P. Reviews. ICES Journal of Marine Science, 1(2):173–174, May 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
2/173/646505.

Thursby-Pelham:1927:BSS

[TP27] D. E. Thursby-Pelham. Biological statistics of the stocks of fish: III. Fluctuations in the age composition of the plaice stock in the Southern North Sea during the period 1923 to 1926. ICES Journal of Marine Science, 2(2):173–183, July 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/2/2/173/650816.

T-P:1928:Ra

[TP28a] D. T.-P. Reviews. ICES Journal of Marine Science, 3(2):264–265, September 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
2/264/624177.

T-P:1928:Rb

[TP28b] D. E. T.-P. Reviews. ICES Journal of Marine Science, 3(3):395–397, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/395/690399.

T-P:1929:R

[TP29] D. E. T.-P. Reviews. ICES Journal of Marine Science, 4(3):354–356, December 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/4/
3/354/653800.

Volterra:1928:VFN

[Vol28] Vito Volterra. Variations and fluctuations of the number of individuals in animal species living together. ICES Journal of Marine Science, 3(1):3–51, April 1928. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/3/1/3/737415.

vanRiel:1929:OIS

[vR29] P. M. van Riel. Oberflächensalzgehalt im südlichen Teil der Nordsee: (Monatskarten, Monats- und Dekaden-Abweichungen). (German) [Surface salt content in the southern part of the North Sea: (monthly maps, monthly and decade deviations)]. ICES Journal of Marine Science, 4 (1):43–67, March 1929. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/4/1/43/744041.

W:1926:R

[W.26] W. Reviews. ICES Journal of Marine Science, 1(1):67–68, February 1926. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/1/
1/67/627649.

Wattenberg:1929:DAO

[Wat29] H. Wattenberg. Die Durchlüftung des Atlantischen Ozeans: (Vorläufige Mitteilung aus den Ergebnissen der Deutschen Atlantischen Expedition). (German) [Ventilation of the Atlantic Ocean: (preliminary information from the results of the German Atlantic Expedition)]. ICES Journal of Marine Science, 4(1):68–79, March 1929. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/4/1/68/744047.

Wimpenny:1928:TBD

[Wim28] M. Sc. R. S. Wimpenny. A transparent bucket with a detachable bottom for use with Ostenfeld and Jespersen’s standard net for plankton collection. ICES Journal of Marine Science, 3(1):94–97, April 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/3/1/94/737425.

Wollaston:1928:NAM

[Wol28] H. J. B. Wollaston. New apparatus for measuring fish. ICES Journal of Marine Science, 3(3):380–381, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/3/3/380/690371.

Yonge:1927:R

[Yon27] C. M. Yonge. Reviews. ICES Journal of Marine Science, 2(1):91–93, March 1927. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/2/
1/91/633989.

Zenkewitch:1928:MSP

[ZBI28] L. Zenkewitch, V. Brotzky, and M. Idelson. Materials for the study of the productivity of the sea-bottom in the White, Barents and Kara Seas. ICES Journal of Marine Science, 3(3):371–379, December 1928. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/3/
3/371/690353.